OpenStudy (anonymous):

Need help factoring 0=3cos^2x-sin^2x

OpenStudy (plasmataco):

welll....

OpenStudy (anonymous):

I got it down to 3cos^2x-1-sin^2x=0

OpenStudy (plasmataco):

wut...

OpenStudy (anonymous):

Oops I meant 3cos^2x-1-cos^2x=0

OpenStudy (anonymous):

I used a trig identity. I changed sin^2x from the problem and changed it to 1-cos^2x

OpenStudy (anonymous):

Anyone here??

OpenStudy (plasmataco):

Im so sry. My internet froze.

OpenStudy (anonymous):

Oh it's ok

jhonyy9 (jhonyy9):

can you continue it - ?

OpenStudy (anonymous):

No I'm stuck on the rest. I know I'm factoring a cos but I'm just confused

jhonyy9 (jhonyy9):

3cos^2x-1-cos^2x=0 what you can doing here - with the same terms ?

OpenStudy (anonymous):

Cancel them out??

jhonyy9 (jhonyy9):

like 3x-1-x = ?

OpenStudy (anonymous):

2x-1??

jhonyy9 (jhonyy9):

so now substitute the cos^2 x in place of x what will get ?

OpenStudy (anonymous):

2cos^2x??

jhonyy9 (jhonyy9):

and with -1 what will be ?

OpenStudy (plasmataco):

Ur almost there

OpenStudy (anonymous):

2cos^2x-1??

jhonyy9 (jhonyy9):

you have forgot it - yes ?

OpenStudy (anonymous):

??

jhonyy9 (jhonyy9):

yes - so than you see it what you see there ? not is like a^2 -b^2 = ?

jhonyy9 (jhonyy9):

formula of difference of two squares - yes ?

OpenStudy (anonymous):

Yes

jhonyy9 (jhonyy9):

so factorize it and finish

jhonyy9 (jhonyy9):

do you know formula sure - yes ?

OpenStudy (anonymous):

No

jhonyy9 (jhonyy9):

Franky what is the problem now here ? or you dont know this formula ?

OpenStudy (anonymous):

I don't know formula

jhonyy9 (jhonyy9):

this will be the last step

OpenStudy (anonymous):

How u do the formula?

jhonyy9 (jhonyy9):

2cos^2 x -1 = ? a^2 - b^2 = ?

OpenStudy (anonymous):

I don't know

jhonyy9 (jhonyy9):

formula for difference of two squares - so this you need to know it

OpenStudy (anonymous):

I don't know

jhonyy9 (jhonyy9):

than i will give you - will remember it sure - i think a^2 -b^2 =(a-b)(a+b)

OpenStudy (anonymous):

So 2cosx-1 and cosx+1 ????

jhonyy9 (jhonyy9):

2 there not is right sure because 2 squared not will give you never 2 - yes ? how many squared will result to you 2 ?

OpenStudy (anonymous):

Why no 2?

jhonyy9 (jhonyy9):

so than you dont believe me try making the proof multiplie what you have wrote above these factors and see will get what you need

jhonyy9 (jhonyy9):

but for give you answer why not 2 so because 2 squared is 4 and not 2

jhonyy9 (jhonyy9):

and than you see on the right part there is 2cos^2 x - and this is the ,,a" - ok ?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

The actual problem says sin^2x=3cos^2x..... I'm solving a trig equation....I subtracted sin^2x on both sides. I then end up with 3cos^2x-sin^2x=0

OpenStudy (anonymous):

\[2 \cos ^{2}2x=1,1+\cos 4x=1,\cos 4x=0=\cos \left( 2n+1 \right)\frac{ \pi }{ 2 },4x=\left( 2n+1 \right)\frac{ \pi }{ 2 }\] x=? n is an integer.

jhonyy9 (jhonyy9):

yes but from the right side you can writing sin^2 x = 1-cos^2 x and so in this way you will get exactly where i have arrived using this formula of difference of two squares - try it than you dont believe me